A closed container of volume 0.02m 3 contains a mixture of neon and argon gases at a temperature 300k and pressure 1*10 5 Nm -2 . The total mass of mixture is 28g . If the molecular weigth of neon and argon are are 20 and 40 resp., find the masses of individual gasae in the container, assuming them to be ideal. PLZ REPLY FAST

A  closed container of volume 0.02m3 contains a mixture of neon and argon gases at a temperature 300k and pressure 1*105 Nm-2 . The total mass of mixture is 28g . If the molecular weigth of neon and argon are are 20 and 40 resp., find the masses of individual gasae in the container, assuming them to be ideal.


 


PLZ REPLY FAST Smile

Grade:11

2 Answers

Chetan Mandayam Nayakar
12 years ago

PV=nRT, n=PV/RT, let weight of neon=x, argon=y, x+y=28, x/20 +y/40 =n=PV/RT, solve the two equations for x and y. 

MOHIT GAUTAM
12 years ago
Thnx :-)